# Supervision, Writing – review and editing

### v1

@@ 4.6.1 Uncertainty increases with prediction distance
Here, we analyze the relationship between prediction distance and uncertainty. For each predicted follow-up MRI, the prediction distance ($\Delta B_{i}$) is calculated as the difference between the follow-up age ($B_{i}$) and the starting age ($A$). The final uncertainty difference ($\Delta u_{i}$) is computed as the difference between the uncertainty at the $i$-th prediction ($u_{i}$) and the uncertainty at the first prediction ($u_{1}$). We use $\Delta B_{i}$ and $\Delta B_{i}^{2}$ as independent variables to predict $\Delta u_{i}$, with subject ID as a random effect. We find that the uncertainty significantly increases with prediction distance $\Delta B_{i}$ ($\beta$ = 0.243, p < 0.001) and $\Delta B_{i}^{2}$ ($\beta$ = 0.460, p < 0.001). Figure 4-A shows the fixed effects from the mixed-effects model, along with the observed values of these variables for individual subjects.
4.6.2 Uncertainty associates with prediction error
@@
Here, we analyze the relationship between the proposed uncertainty measure and image-based metrics (MSE and SSIM). Specifically, we use the square of the uncertainty ($u^{2}$) as the independent variable to predict the image-based metrics, while incorporating subject ID as a random effect. Our results demonstrate a significant positive correlation between MSE and $u^{2}$ ($\beta=0.157$, p < 0.001), indicating that higher uncertainty corresponds to increased mean squared error. At the same time, we find that SSIM exhibits a significant negative correlation with $u^{2}$ ($\beta=-2.008$, p < 0.001), suggesting that higher uncertainty is associated with decreased structural similarity. We illustrate these relationships in Figure 4-B and Figure 4-C, which present both the fixed effects estimated by our model and the observed values from individual subjects. These findings collectively support the utility of our uncertainty measure as a predictor of image quality metrics in the context of BrLP predictions.
@@ 4.7 Downstream application: avoiding type II errors in clinical trials using BrLP
Type II errors in clinical trials occur when a study fails to demonstrate the efficacy of a treatment that actually has significant effects. This can result in substantial financial losses and the failure to bring potentially beneficial drugs to market. One contributing factor to Type II errors is the inclusion of patients whose disease progression is too slow to observe treatment effects within the trial duration. Consequently, there is growing interest in identifying and selecting "fast progressors" - subjects whose disease advances more rapidly. Predictive tools such as BrLP can be used for this downstream application, providing predictive capabilities to assess disease progression at an individual level and facilitating the identification of fast progressors.
@@
Figure 5 illustrates the results of these experiments. We report results as improvements over the random selection strategy, averaged across all $S$. The regression model achieves the best performance, selecting 25.65% more fast progressors than random selection in the internal test set and 18.83% more in the external test set (representing a 6.82% drop in performance). BrLP produces comparable results, identifying 20.45% more fast progressors than random selection in the internal test set and 15.28% more in the external dataset (a 5.17% drop in performance). Although less accurate than the regression-based strategy, BrLP shows better consistency across different datasets, exhibiting a smaller performance drop on the external dataset compared to the regression model. This suggests that BrLP may be more robust and generalizable to changes in the dataset. These results underscore BrLP’s potential as a tool for patient selection in clinical trials. Its ability to better generalize across datasets highlights its suitability for downstream tasks in clinical research, even though it is not specifically optimized for this purpose.
@@ 5 Discussion
In Section 4.5, we demonstrate that BrLP effectively exploits the conditioning covariates, yielding better quantitative results when using the correct cognitive status of each test subject. In Section 4.3, we show that incorporating the auxiliary model enhances BrLP’s performance, highlighting the importance of exploiting longitudinal data when developing spatiotemporal models. Additionally, we illustrate in the same section that implementing the LAS algorithm improves overall performance, though it introduces a trade-off between computational overhead—due to the hyperparameter $m$—and prediction accuracy. Section 4.6 demonstrates that the uncertainty measure derived from the LAS algorithm is associated with both prediction distance and prediction error, offering a potential tool for assessing prediction reliability in clinical applications. Finally, in Section 4.7, we showcase how BrLP predictions can identify patients with faster disease progression during clinical trials, potentially reducing the risk of type II errors.
While these advancements underscore BrLP’s strengths, in our experiments, we also identify a key limitation of BrLP: a bias in its predictive performance, with quantitative metrics generally being less accurate for AD cases compared to healthy aging (see Section 4.4). This discrepancy is likely due to the inherent heterogeneity and complexity of AD progression, which pose significant challenges for reliable prediction. Interestingly, a similar bias is observed in all the other baseline models, suggesting that this is not unique to BrLP and highlighting a broader avenue for future research.
Future research will aim to reduce this bias by further disentangling disease progression from the ageing process and incorporating additional disease-specific variables that influence the severity and pace of progression in individuals. One potential direction is to model disease stage transitions (e.g., from MCI to AD) during longitudinal predictions, rather than assuming the subject’s final cognitive status. Furthermore, extending uncertainty estimation to the voxel level will also be critical, offering more precise visualizations of model confidence across different regions of the predicted brain MRI. Finally, adapting and applying BrLP to other brain diseases and chronic conditions affecting various organs will further demonstrate its versatility and clinical impact.

### v2

@@ 5.6.1 Global-level uncertainty increases with prediction distance
Here, we analyze the relationship between prediction distance and global-level uncertainty. For each predicted follow-up MRI, the prediction distance ($\Delta B_{i}$) is calculated as the difference between the follow-up age ($B_{i}$) and the starting age ($A$). The final uncertainty difference ($\Delta u_{i}$) is computed as the difference between the uncertainty at the $i$-th prediction ($u_{i}$) and the uncertainty at the first prediction ($u_{1}$). We use $\Delta B_{i}$ and $\Delta B_{i}^{2}$ as independent variables to predict $\Delta u_{i}$, with subject ID as a random effect. We find that the global-level uncertainty significantly increases with prediction distance $\Delta B_{i}$ ($\beta$ = 0.243, p < 0.001) and $\Delta B_{i}^{2}$ ($\beta$ = 0.460, p < 0.001). Figure 5-A shows the fixed effects from the mixed-effects model, along with the observed values of these variables for individual subjects.
5.6.2 Global-level uncertainty associates with prediction error
@@
Here, we analyze the relationship between the proposed global-level uncertainty measure and image-based metrics (MSE and SSIM). Specifically, we use the square of the uncertainty ($u^{2}$) as the independent variable to predict the image-based metrics, while incorporating subject ID as a random effect. Our results demonstrate a significant positive correlation between MSE and $u^{2}$ ($\beta=0.157$, p < 0.001), indicating that higher uncertainty corresponds to increased mean squared error. At the same time, we find that SSIM exhibits a significant negative correlation with $u^{2}$ ($\beta=-2.008$, p < 0.001), suggesting that higher global-level uncertainty is associated with decreased structural similarity. We illustrate these relationships in Figure 5-B and Figure 5-C, which present both the fixed effects estimated by our model and the observed values from individual subjects. These findings collectively support the utility of our global-level uncertainty measure as a predictor of image quality metrics in the context of BrLP predictions.
5.6.3 Evaluating the voxel-level uncertainty map
In this section, we evaluate the voxel-level uncertainty map using both quantitative and qualitative analyses. For the quantitative assessment, we adopt the methodology of (Zhou et al., 2024) and compute the correlation between the uncertainty map $U^{(B)}$ and voxel-wise squared prediction error (between predicted and ground-truth follow-up images). We explicitly exclude background voxels to prevent artificial inflation of the correlation. We obtain a Spearman correlation coefficient of $0.63\pm 0.11$, suggesting that the model assigns higher uncertainty to regions with greater prediction error. In Figure 6, we showcase a qualitative example of the obtained voxel-level map, illustrating how uncertainty evolves with increasing prediction distance. We observe a gradual rise in uncertainty around the lateral ventricles, consistent with the findings reported in Section 5.6.1. Moreover, the uncertainty map visually corresponds with the error map in the latest predictions, reinforcing the quantitative results. Together, these findings suggest that the model’s uncertainty estimates are both meaningful and informative for assessing prediction reliability at the voxel level.
@@ 5.7 Downstream application: avoiding type II errors in clinical trials using BrLP
Type II errors in clinical trials occur when a study fails to demonstrate the efficacy of a treatment that actually has significant effects. This can result in substantial financial losses and the failure to bring potentially beneficial drugs to market. One contributing factor to Type II errors is the inclusion of patients whose disease progression is too slow to observe treatment effects within the trial duration. Consequently, there is growing interest in identifying and selecting "fast progressors" - subjects whose disease advances more rapidly (Wang et al., 2024; Birkenbihl et al., 2024). Predictive tools such as BrLP can be used for this downstream application, providing predictive capabilities to assess disease progression at an individual level and facilitating the identification of fast progressors.
@@
Figure 7 presents the results of these experiments. We report performance as improvements over a random selection strategy, averaged across all values of $S$. The regression model achieves the highest performance, selecting 25.67% more fast progressors than random selection on the internal test set, and 24.83% more on the external test (representing a 0.84% drop in performance). BrLP achieves similar results, identifying 24.94% more fast progressors than random selection internally and 23.85% externally (a 1.09% drop in performance). Although BrLP was not explicitly trained for this task, its performance is comparable to that of the regression model, highlighting its potential for downstream clinical applications such as patient selection for clinical trials.
@@ 6 Discussion
In Section 5.5, we demonstrate that BrLP effectively exploits the conditioning covariates, yielding better quantitative results when using the correct cognitive status of each test subject. In Section 5.3, we show that incorporating the auxiliary model enhances BrLP’s performance, highlighting the importance of exploiting longitudinal data when developing spatiotemporal models. Additionally, we illustrate in the same section that implementing the LAS algorithm improves overall performance, though it introduces a trade-off between computational overhead—due to the hyperparameter $m$—and prediction accuracy. Section 5.6 demonstrates that the global- and voxel-level uncertainty measures derived from the LAS algorithm are associated with both prediction distance and prediction error, offering a potential tool for assessing prediction reliability in clinical applications. Finally, in Section 5.7, we showcase how BrLP predictions can identify patients with faster disease progression during clinical trials, potentially reducing the risk of type II errors.
While our results highlight BrLP’s strengths, some experiments also reveal a few limitations. The first limitation is a bias in predictive performance: quantitative metrics are generally less accurate for AD cases compared to healthy aging (see Section 5.4). This discrepancy is likely due to the inherent heterogeneity and complexity of AD progression. Interestingly, a similar bias is observed across all baseline models, suggesting that it is not unique to BrLP. The second limitation involves a possible sex-related bias: as shown in Appendix D, the model performs slightly better on female subjects in terms of image-based metrics. A third limitation, as also highlighted in (McMaster et al., 2025), is a smoothing effect in BrLP’s outputs. This effect is likely introduced by the VAE component (Rivera, 2024) and may reduce the model’s ability to capture fine-grained details in the data.
Future research will aim to reduce the performance gap between the healthy and AD predictions by further disentangling disease progression from the ageing process and incorporating additional disease-specific variables that influence the severity and pace of progression in individuals. One potential direction is to model disease stage transitions (e.g., from MCI to AD) during longitudinal predictions, rather than assuming the subject’s final cognitive status. Finally, adapting and applying BrLP to other brain diseases and chronic conditions affecting various organs will further demonstrate its versatility and clinical impact.
@@ A.2 Disease Course Mapping (DCM)
Appendix B Dataset statements
(i) Data was collected by the AIBL study group. AIBL study methodology has been reported previously (Ellis et al., 2009) . (ii) Data used in the preparation of this article were obtained from the Alzheimer’s Disease Neuroimaging Initiative (ADNI) database (adni.loni.usc.edu). The ADNI was launched in 2003 as a public-private partnership, led by Principal Investigator Michael W. Weiner, MD. The original goal of ADNI was to test whether serial MRI, positron emission tomography (PET), other biological markers, and clinical and neuropsychological assessment can be combined to measure the progression of MCI and early AD. The current goals include validating biomarkers for clinical trials, improving the generalizability of ADNI data by increasing diversity in the participant cohort, and to provide data concerning the diagnosis and progression of AD to the scientific community. For up-to-date information, see adni.loni.usc.edu.
Figure 8: Effect of the number of DDIM inference steps on BrLP performance. SSIM (left axis, blue) and MSE (right axis, red) are reported for different numbers of denoising steps. Shaded areas indicate 95% confidence intervals.
Appendix C Impact of reducing DDIM denoising steps on BrLP’s performance
We conduct an ablation study to evaluate how lowering the number of DDIM (Song et al., 2021) denoising steps affects BrLP’s performance. Specifically, we compare the image-based metrics obtained using 25 steps (as in our experiments) to those obtained with 15, 10, 5, and 2 steps. The results, shown in Figure 8, reveal that reducing the number of steps to as few as 5 maintains satisfactory performance, with only a minor decline in quality.
Table 5: Evaluation of BrLP performance differences between male and female subjects. MSE and regional MAE values (± SD) are reported, with the best result for each metric between the two groups highlighted in bold. A result is marked with a star if it is significantly better than all other configurations at the 5% significance level (Mann–Whitney U test).
Internal
test set
Males Single image 0.005 ± 0.003 0.909 ± 0.032 0.021 ± 0.019* 0.016 ± 0.014 0.273 ± 0.303 0.029 ± 0.023 0.828 ± 0.630
Females 0.004 ± 0.002* 0.919 ± 0.022* 0.024 ± 0.022 0.014 ± 0.013 0.222 ± 0.294* 0.028 ± 0.022 0.823 ± 0.616
[…]
Females 0.004 ± 0.002* 0.918 ± 0.021* 0.020 ± 0.016 0.013 ± 0.012 0.199 ± 0.218* 0.029 ± 0.024 0.773 ± 0.577
External
test set
Males Single image 0.005 ± 0.002 0.906 ± 0.023 0.025 ± 0.021 0.015 ± 0.013 0.210 ± 0.240 0.030 ± 0.023 1.036 ± 0.736
Females 0.004 ± 0.002* 0.910 ± 0.023* 0.024 ± 0.024 0.013 ± 0.013 0.215 ± 0.402 0.030 ± 0.024 1.049 ± 0.818
[…]
Females 0.004 ± 0.002* 0.913 ± 0.021* 0.022 ± 0.023 0.012 ± 0.012* 0.207 ± 0.430 0.030 ± 0.023 1.000 ± 0.775
Appendix D Analysis of sex differences in BrLP’s predictions
To assess potential sex-related differences in BrLP’s performance, we stratify the evaluation by sex on both internal and external test sets, under single-image and sequence-aware settings. The results are summarized in Table 5. BrLP demonstrates slightly better performance for female subjects on image-based metrics in both test sets (Mann–Whitney U test, $p<0.05$). Furthermore, volumetric analysis of the lateral ventricles shows a significantly lower MAE for female subjects in the internal test set ($p<0.05$), although this difference does not persist in the external test set. In conclusion, we find that BrLP exhibits limited sex-related bias across both image-level and volumetric evaluations.